A fabrication company engaged in production with a capacity of 150, 000 pieces per year. But, it is just operating at 70% of its full capacity. The company has an annual income of 250, 000, annual fixed cost are 50, 000 and variable costs are 1.00 pesos per unit. How many productions of parts must be produced for break-even point?


1
Expert's answer
2022-01-20T09:54:36-0500

The price is:

"P = (TP + TC)\/Q = (250,000 + 50,000 + 1\u00d70.7\u00d7150,000)\/(0.7\u00d7150,000) = 3.86."

The break-even point is:

"Q = 50,000\/(3.86 - 1) = 17,482.5" units.